OpenStudy (anonymous):

the indefinite integral i am given is (e^x - 6x^2)/6... i need to evaluate it. please help!

OpenStudy (nowhereman):

Do you know the integral of e^x and the power rule?

OpenStudy (anonymous):

i assume the integral is e^x

OpenStudy (anonymous):

i need to use substitution, but i dont know what to substitute... i've tried it a lot already

OpenStudy (nowhereman):

You don't need to use any elevated integration techniques. If you want to prove that your result is right, just differentiate it and if you get the original integrand, you're fine.

OpenStudy (anonymous):

what about the 6 in the denominator? how do i do that? is it just 6x?

OpenStudy (nowhereman):

No, that is just a factor of 1/6 which you can pull-out.

OpenStudy (anonymous):

wow.. i got it! thanks! i have another one tho if u are willing to help?

OpenStudy (nowhereman):

Sure :-)

OpenStudy (anonymous):

(5-2xe^x)/x

OpenStudy (nowhereman):

Pull that division into the difference and you should have two known integrals.

OpenStudy (anonymous):

like (5/x)-((2xe^x)/x)?

OpenStudy (nowhereman):

Yes, and reduce that x in the second term.

OpenStudy (anonymous):

ok. then i would need substitution right?

OpenStudy (nowhereman):

For what? You should know the integral of 1/x and e^x and that's all there is.

OpenStudy (anonymous):

i dont know the 1/x one. is it natural log?

OpenStudy (nowhereman):

Exactly, sometimes it is defined this way \[\ln x := \int_1^x\frac 1 y \mathrm d y\] but if you have defined it as the inverse function of e^x you could use the inverse-differentiation rule.

OpenStudy (anonymous):

i got that one! thanks a lot. this is my last one... (x-2)/(x^2-4x+4)^3..... that requires substitution correct?

OpenStudy (nowhereman):

Yes, a little one :-) but you should first observe that \[x^2 - 4x + 4 = (x-2)^2\]

OpenStudy (anonymous):

so i can cancel one of them right away?

OpenStudy (nowhereman):

yes

OpenStudy (anonymous):

would it turn into (x-2)^6?

OpenStudy (anonymous):

in the denominator

OpenStudy (nowhereman):

Sure, that is just one of the exponentiation rules.

OpenStudy (anonymous):

then after the cancellation i am left with 1/(x-2)^5?

OpenStudy (nowhereman):

Right.

OpenStudy (anonymous):

now i substitute u=x-2

OpenStudy (nowhereman):

Yes, but as du/dx = 1 if you are a little more experienced you could just do it in your head.

OpenStudy (anonymous):

i think it would be 1/((1/6)*(x-2)^6)

OpenStudy (nowhereman):

Not quite, maybe you should write it is \[(x-2)^{-5}\] instead to see it better.

OpenStudy (anonymous):

(-1/4)*(x-2)^-4?

OpenStudy (nowhereman):

Yeah, good.

OpenStudy (anonymous):

is this website free to use?

OpenStudy (nowhereman):

It costs both of us our time, but you won't have to pay for it ;-)

OpenStudy (anonymous):

so y do ppl like u go around and answer questions?

OpenStudy (nowhereman):

I love math and helping other people.

OpenStudy (anonymous):

it seems strange that there is no charge for this service. or breech of security

OpenStudy (nowhereman):

I don't know what you mean by breech of security; but the idea of being is just that everyone can ask questions and give answers here. So if I had a question regarding my study, I could ask them too. Only that it seems there are a lot of high-school students here lately :-D

OpenStudy (anonymous):

hmm... thats cool. i guess i'm a skeptic. but with quick responses like the ones u gave me i am becoming a believer